# FINITE SUBGROUPS OF MAXIMAL ORDER OF THE CREMONA GROUP OVER THE RATIONALS

### v1

@@ 7. Del Pezzo surfaces of degree $4$
A pair $\{A,B\}$ of $\mathbb{C}-$divisors of $S$ is an exceptional pair if $A$ and $B$ are $\mathbb{C}-$divisors, $A^{2}=B^{2}=0$ and $(A+B)=-K_{S}$. This implies that $AB=2$ and $AK_{S}=BK_{S}=-2$.
@@
    The homomorphism $\gamma:G_{S}\rightarrow(\mathbb{F}_{2})^{5}$, induces an isomorphism $G_{S}\rightarrow(\mathbb{F}_{2})^{4}$.
@@
[…]
@@
    Let $H_{S}$ contains a $3-$cycle, namely $(A_{1}A_{2}A_{5})$. The automorphism is then $\varphi:[x:y:z]\mapsto[a^{2}y:-b^{2}x+aby:acy-abz]$ for some $a,b,c$ in $\mathbb{Q}$ with $a\cdot b\neq 0$. Since $\varphi$ sends $A_{1}$ on $A_{2}$, $A_{2}$ on $A_{5}$, and $A_{5}$ on $A_{1}$, then $A_{3}$ and $A_{4}$ must be fixed, so we have the equation $\varphi(A_{4})=[a^{2}:-b(a-b):a(c-b)]$, so $a\in\{\omega b,\omega^{2}b\}$, where $\omega$ is the $3rd$ root of unity which is a contradiction.
@@
    Let $H_{S}$ contains a $2-$cycle, namely $(A_{1}A_{5})$. The automorphism is then has to fix the other three points $A_{2},A_{3}$, and $A_{4}$. So it is given by $\varphi:[x:y:z]\mapsto[ax:bx+(a-b)y:cx+(a-c)z]$ for some $a,b,c$ in $\mathbb{Q}$ with $a\cdot b\cdot c\neq 0$. Since $\varphi$ sends $A_{1}$ on $A_{5}$, then $A_{5}$ has to be sent on $A_{1}$, but $\varphi(A_{5})=[a^{2}:2ab-b^{2}:2ac-c^{2}]=[1:0:0]$, so we have either $c=b$ and $a=b/2$ which is a contradiction: hence $A_{5}=[1,2,2]$ which is collinear with $A_{1}$ and $A_{4}$.
@@
    Let $H_{S}$ contains a $5-$cycle, namely $(A_{1}A_{4}A_{3}A_{2}A_{5})$. The automorphism is then $[x:y:z]\mapsto[a(x-y):ax-by-(a-b)z:ax-cy]$ for some $a,b,c$ in $\mathbb{Q}$ with $a\cdot b\neq 0$. Since the automorphism sends $A_{1}$ on $A_{4}$, $A_{4}$ on $A_{3}$, $A_{3}$ on $A_{2}$, and $A_{2}$ on $A_{5}$, then the image of $A_{5}$ is $[a^{2}-ab:a^{2}-b^{2}-(a-b)c:a^{2}-bc]$ must be $A_{1}$, then $a^{3}-2a^{2}b-b^{3}=0$, assuming $a=t\cdot b$ implies that $b^{3}(t^{3}-2t^{2}-1)=0$. then either $a=b=0$ or $t^{3}-2t^{2}-1=0$, but the discriminant of the polynomial $t^{3}-2t^{2}-1$ is $-59$, then polynomial $t^{3}-2t^{2}-1$ has one real root and two complex conjugate roots, and then the real root is $x_{0}=(r^{\frac{2}{3}}+4r^{\frac{1}{3}}+16)/(6r^{\frac{1}{3}})$ where $r=172+12\sqrt{177}$, which is not rational, so there is no non-trivial solution in $\mathbb{Q}$, which is a contradiction.
As $H_{S}$ only contains $2\times 2-$cycle, it has order $1,~{}2$ or $4$. One then needs to remove the case where $\mid H_{S}\mid=4$ to finish the proof. In that case, $H_{S}$ would be, up to conjugation, the Klein group $\langle(A_{1}A_{2})(A_{3}A_{4}),(A_{1}A_{3})(A_{2}A_{4})\rangle$ and thus would be generated by $[x:y:z]\mapsto[z-y:z-x:z]$ and $[x:y:z]\mapsto[y-z:y:y-x]$, so $A_{5}$ should be fixed by both elements. However, the only three points of $\mathbb{P}^{2}(\mathbb{C})$ fixed by these two involutions are [0:1:1],[1:0:1][0:1:1],[1:0:1][ 0 : 1 : 1 ] , [ 1 : 0 : 1 ] and $[1:1:0]$, and any of these is collinear with two of the points $A_{1},A_{2},A_{3},A_{4}$. This leads to a contradiction, that shows that $H_{S}$ has order $1$ or $2$, so $\operatorname{Aut}(S)$ has order $16$ or $32$. ∎
@@
This gives to $X$ a unique structure of a surface defined over $\mathbb{Q}$. It contains the $\mathbb{Q}$-rational point ([1:1:1],[1:1:1])([1:1:1],[1:1:1])( [ 1 : 1 : 1 ] , [ 1 : 1 : 1 ] ) and is thus rational [Man66, Theorem 3.15]. We choose the two other $\mathbb{Q}$-rational points
@@
denote by $Y$ the blow-up $Y$ of $X$ at the two points $q_{1}$, $q_{2}$. We now prove that $\operatorname{Aut}(Y)$ is isomorphic to $(\mathbb{Z}/2\mathbb{Z})^{4}\rtimes\mathrm{Sym}_{3}$ and thus has $962^{4}\cdot 6=$ 96 elements.
@@
Let us assume that $S$ is a Del Pezzo surface of degree $4$ defined over $\mathbb{Q}$. Let $L/\mathbb{Q}$ be the smallest Galois field extension such that the $(-1)-$curves on $S$ are defined over $L$. The surface $S$ has sixteen $(-1)-$curves defined over $L$ but not necessarily over $\mathbb{Q}$. By Lemma 7.3, we have $\operatorname{Aut}_{\mathbb{C}}(S)\cong G_{S}\rtimes H_{S}$, $G_{S}\cong(\mathbb{F}_{2})^{4}$, and the group $H_{S}\subset\operatorname{Sym}_{5}$ is isomorphic to either $\mathbb{Z}/2\mathbb{Z},~{}\mathbb{Z}/4\mathbb{Z},~{}\operatorname{Sym}_{3}$ or $D_{5}$ (the dihedral group of order $10$). Since the surface $S$ is defined over $\mathbb{Q}$, then $\operatorname{Aut}_{\mathbb{Q}}(S)\subseteq\operatorname{Aut}_{L}(S)\subseteq\operatorname{Aut}_{\mathbb{C}}(S)$. If $\mid\operatorname{Aut}_{\mathbb{Q}}(S)\mid>96$, then $\operatorname{Aut}_{\mathbb{Q}}(S)\cong(\mathbb{F}_{2})^{4}\rtimes D_{5}$. But in this case, every element in the Galois group $\operatorname{Gal}(L/\mathbb{Q})$ should commute with the whole group $D_{5}$. Since the only element of the group $(\mathbb{F}_{2}^{4})\rtimes D_{5}$ that commutes with that whole group is the identity, then the Galois group acts trivially. It follows that every $(-1)-$curve is defined over $\mathbb{Q}$, hence by Lemma 7.4, we know that $\mid\operatorname{Aut}_{\mathbb{Q}}(S)\mid\leq 32$ which is a contradiction, so $H_{S}$ can not be $D_{5}$. ∎
8. Del Pezzo surfaces of degree $3$
Let $S$ be a rational Del Pezzo surface of degree $3$ defined over $\mathbb{Q}$, we want to find the biggest automorphism group $G\subseteq\operatorname{Aut}_{\mathbb{Q}}(S)$. The classification of all automorphism groups of $S$ over $\mathbb{C}$ has been studied in [Dol12, Theorem 9.5.6], and they explained that the Del Pezzo surface of degree $3$ is isomorphic to a one contained in one of the $11$ families of smooth cubic surfaces given in the list.
@@ 8. Del Pezzo surfaces of degree $3$
In Theorem 6.1, since $M(4,2)=7,~{}M(4,3)=2,~{}M(4,5)=1,$ and $M(4,p)=0$ for all primes $p>5$, then we have
@@
Let $G$ be a finite subgroup of $\operatorname{PGL}_{n}(\mathbb{Q})$ and let $F\in\mathbb{Q}[x_{0},\ldots,x_{n-1}]$ be an irreducible homogeneous polynomial such that $G$ preserves $F$. This means that for each $g$ in $G$ and each matrix $M$ of $\mathrm{GL}_{n}(\mathbb{Q})$ such that g is the class of $M$, the polynomial $F\circ M$ (obtained by acting on the coordinates linearly) is a multiple of $F$. Then, if $d=deg(F)$ and $n$ are coprime, there exists a finite subgroup of $\mathrm{GL}_{n}(\mathbb{Q})$ that surjects to $G$, is twice bigger, and contains $-Id$.
@@
Let $g$ be any element of $G$ and let $M$ be the corresponding matrix. We have $F\circ M=\lambda\cdot F$ for some $\lambda\in\mathbb{Q}$. Let $m$ be the order of $g$. Then, $M^{m}$ is a multiple of the identity, which gives $\mu\in\mathbb{Q}$ such that $M^{m}=\mu\cdot Id$. We write $a=det(M)$. As $M^{m}=\mu\cdot Id$, we apply the determinant on both sides and obtain $a^{m}=\mu^{n}$. Moreover, $F\circ M=\lambda\cdot F$, so $F\circ(M^{i})=\lambda^{i}\cdot F$ for each $i\geq 1$. In particular, $F\circ(M^{m})=\lambda^{m}\cdot F$. As $M^{m}=\mu\cdot Id$, we find $F\circ(\mu\cdot Id)=F\cdot\mu^{d}$ (recall that $d$ is the degree of $F$), so $\mu^{d}=\lambda^{m}$. By Lemma 6.3, there exist $c\in\mathbb{Q}^{*}$ such that $\mu=c^{m}$. Replacing $M$ with $M/c$ we do not change the class $g$ in $G$, but may assume that $M^{m}=Id$. This implies that $\mathrm{det}(M)^{m}=1$, so $\mathrm{det}(M)\in\{-1,1\}$. We can then lift any element of $g\in G$ to an element of $\mathrm{GL}_{n}(\mathbb{Q})$ that has determinant $1$ or $-1$. This gives exactly two possibilities for the lift, namely $M$ or $-M$. The set of such matrices makes a finite group twice bigger than $G$. ∎
@@
Let $S$ be a rational Del Pezzo surface defined over $\mathbb{Q}$. Since $\operatorname{Aut}_{\mathbb{Q}}(S)\subseteq\operatorname{Aut}_{\mathbb{C}}(S)$, by [Dol12, Theorem 9.5.6], we only need to study the automorphism groups $G\subset\operatorname{Aut}_{\mathbb{C}}(S)$ such that $\mid G\mid>120$, hence we have only one type to study. If the surface is isomorphic to the Fermat surface $V(t_{0}^{3}+t_{1}^{3}+t_{2}^{3}+t_{3}^{3})$, then $\operatorname{Aut}_{\mathbb{C}}(S)\cong(\mathbb{Z}/3\mathbb{Z})^{3}\times\mathfrak{S}_{4}$ where $\mathfrak{S}_{4}$ is the group of permutations, then $\mid G\mid$ has to divide $648=2^{3}\cdot 3^{4}$. Since $G\subseteq\operatorname{PGL}_{4}(\mathbb{Q})$, then by Remark 8.3, $\mid G\mid$ has to divide $2^{6}\cdot 3^{2}\cdot 5$, hence $\mid G\mid\leq 2^{3}\cdot 3^{2}=48$. ∎
@@
This surface $X$ contains $2$ disjoint $(-1)-$curves defined over $\mathbb{Q}$ , given by
$\begin{array}[]{l}E_{1}=\{[0:a:-a:b:-b]\mid[a:b]\in\mathbb{P}^{1}\},\\ E_{2}=\{[a:0:b:-a:-b]\mid[a:b]\in\mathbb{P}^{1}\}.\\ \end{array}$
By contracting these two $(-1)-$curves, there is thus a contraction morphism $\eta:X\rightarrow S_{5}$, where $S_{5}$ is a Del Pezzo surface of degree $5$ with at least $2$ rational points, so the surface $S_{5}$ is a rational surface. Indeed, this follows from [Man66, Theorem 3.15]. Furthermore, the group $G\cong\mathfrak{S}_{5}$ has order $120$.

### v2

@@ 7. Del Pezzo surfaces of degree 44
A pair $\{A,B\}$ of classes of $\mathbb{C}$-divisors where $A,B\in Pic_{C}(S)$ is an exceptional pair if $A$ and $B$ are $\mathbb{C}$-divisors, $A^{2}=B^{2}=0$ and $(A+B)=-K_{S}$. This implies that $AB=2$ and $AK_{S}=BK_{S}=-2$.
@@
    The homomorphism $\gamma:G_{S}\rightarrow(\mathbb{F}_{2})^{5}$, induces an isomorphism $G_{S}\rightarrow\{(a_{0},a_{1},a_{2},a_{3},a_{4})\in(\mathbb{F}_{2})^{5}\mid\Sigma_{i=0}^{4}a_{i}=0\}\cong(\mathbb{F}_{2})^{4}$.
@@
    Let $H_{S}$ contains a 44-cycle, namely $(A_{1}A_{3}A_{2}A_{4})$. The automorphism is then $x:y:z]\mapsto[y:y-z:y-x]$. The point $A_{5}$ must be a fixed point, so we have either $b=\mathrm{i}a$ where $\mathrm{i}=\sqrt{-1}$ or $\{b=a,c=0\}$, then we have a contradiction: it is clear in the first case as $b\notin\mathbb{Q}$ and in the second case, we have $[a:b:c]=[1:1:0]$ which is collinear with $A_{1}$ and $A_{2}$.
@@
    Let $H_{S}$ contains a 33-cycle, namely $(A_{1}A_{2}A_{5})$. The automorphism is then $\varphi:[x:y:z]\mapsto[a^{2}y:-b^{2}x+aby:acy-abz]$ for some $a,b,c$ in $\mathbb{Q}$ with $a\cdot b\neq 0$. Since $\varphi$ sends $A_{1}$ on $A_{2}$, $A_{2}$ on $A_{5}$, and $A_{5}$ on $A_{1}$, then $A_{3}$ and $A_{4}$ must be fixed, so we have the equation $\varphi(A_{4})=[a^{2}:-b(a-b):a(c-b)]$, so $a\in\{\omega b,\omega^{2}b\}$, where $\omega$ is the $3rd$ root of unity which is a contradiction.
@@
    Let $H_{S}$ contains a 22-cycle, namely $(A_{1}A_{5})$. The automorphism is then has to fix the other three points $A_{2},A_{3}$, and $A_{4}$. So it is given by $\varphi:[x:y:z]\mapsto[ax:bx+(a-b)y:cx+(a-c)z]$ for some $a,b,c$ in $\mathbb{Q}$ with $a\cdot b\cdot c\neq 0$. Since $\varphi$ sends $A_{1}$ on $A_{5}$, then $A_{5}$ has to be sent on $A_{1}$, but $\varphi(A_{5})=[a^{2}:2ab-b^{2}:2ac-c^{2}]=[1:0:0]$, so we have either $c=b$ and $a=b/2$ which is a contradiction: hence $A_{5}=[1,2,2]$ which is collinear with $A_{1}$ and $A_{4}$.
@@
    Let $H_{S}$ contains a 55-cycle, namely $(A_{1}A_{4}A_{3}A_{2}A_{5})$. The automorphism is then $x:y:z]\mapsto[a(x-y):ax-by-(a-b)z:ax-cy]$ for some $a,b,c$ in $\mathbb{Q}$ with $a\cdot b\neq 0$. Since the automorphism sends $A_{1}$ on $A_{4}$, $A_{4}$ on $A_{3}$, $A_{3}$ on $A_{2}$, and $A_{2}$ on $A_{5}$, then the image of $A_{5}$ is $a^{2}-ab:a^{2}-b^{2}-(a-b)c:a^{2}-bc]$ must be $A_{1}$, then $a^{3}-2a^{2}b-b^{3}=0$, assuming $a=t\cdot b$ implies that $b^{3}(t^{3}-2t^{2}-1)=0$. then either $a=b=0$ or $t^{3}-2t^{2}-1=0$, but the discriminant of the polynomial $t^{3}-2t^{2}-1$ is 59-59, then polynomial $t^{3}-2t^{2}-1$ has one real root and two complex conjugate roots, and then the real root is $x_{0}=(r^{\frac{2}{3}}+4r^{\frac{1}{3}}+16)/(6r^{\frac{1}{3}})$ where $r=172+12\sqrt{177}$, which is not rational, so there is no non-trivial solution in $\mathbb{Q}$, which is a contradiction.
As $H_{S}$ only contains $2\times 2$-cycle, it has order 1,21,~2 or $4$. One then needs to remove the case where $\mid H_{S}\mid=4$ to finish the proof. In that case, $H_{S}$ would be, up to conjugation, the Klein group $\langle(A_{1}A_{2})(A_{3}A_{4}),(A_{1}A_{3})(A_{2}A_{4})\rangle$ and thus would be generated by $x:y:z]\mapsto[z-y:z-x:z]$ and $x:y:z]\mapsto[y-z:y:y-x]$, so $A_{5}$ should be fixed by both elements. However, the only three points of $\mathbb{P}^{2}(\mathbb{C})$ fixed by these two involutions are $[0:1:1],[1:0:1]$ and $[1:1:0]$, and any of these is collinear with two of the points $A_{1},A_{2},A_{3},A_{4}$. This leads to a contradiction, that shows that $H_{S}$ has order $1$ or $2$, so $\operatorname{Aut}(S)$ has order $16$ or $32$. ∎
@@
This gives to $X$ a unique structure of a surface defined over $\mathbb{Q}$. It contains the $\mathbb{Q}$-rational point $([1:1:1],[1:1:1])$ and is thus rational [17, Theorem 3.15]. We choose the two other $L$-rational points
@@
denote by $Y$ the blow-up $Y$ of $X$ at the two points $q_{1},~q_{2}$. We now prove that $\operatorname{Aut}(Y)$ is isomorphic to $(\mathbb{Z}/2\mathbb{Z})^{4}\rtimes\mathrm{Sym}_{3}$ and thus has $2^{4}\cdot 6=96$ elements.
@@
Let us assume that $S$ is a Del Pezzo surface of degree $4$ defined over $\mathbb{Q}$. Let $L/\mathbb{Q}$ be the smallest Galois field extension such that the (1)(-1)-curves on $S$ are defined over $L$. The surface $S$ has sixteen (1)(-1)-curves defined over $L$ but not necessarily over $\mathbb{Q}$. By Lemma 7.3, we have $\operatorname{Aut}_{\mathbb{C}}(S)\cong G_{S}\rtimes H_{S}$, $G_{S}\cong(\mathbb{F}_{2})^{4}$, and the group $H_{S}\subset\operatorname{Sym}_{5}$ is isomorphic to either $\mathbb{Z}/2\mathbb{Z},~\mathbb{Z}/4\mathbb{Z},~\operatorname{Sym}_{3}$ or $D_{5}$ (the dihedral group of order $10$). Since the surface $S$ is defined over $\mathbb{Q}$, then $\operatorname{Aut}_{\mathbb{Q}}(S)\subseteq\operatorname{Aut}_{L}(S)\subseteq\operatorname{Aut}_{\mathbb{C}}(S)$. If $\mid\operatorname{Aut}_{\mathbb{Q}}(S)\mid>96$, then $\operatorname{Aut}_{\mathbb{Q}}(S)\cong(\mathbb{F}_{2})^{4}\rtimes D_{5}$. But in this case, every element in the Galois group $\operatorname{Gal}(L/\mathbb{Q})$ should commute with the whole group $D_{5}$. Since the only element of the group $(\mathbb{F}_{2}^{4})\rtimes D_{5}$ that commutes with that whole group is the identity, then the Galois group acts trivially. It follows that every (1)(-1)-curve is defined over $\mathbb{Q}$, hence by Lemma 7.4, we know that $\mid\operatorname{Aut}_{\mathbb{Q}}(S)\mid\leq 32$ which is a contradiction, so $H_{S}$ can not be $D_{5}$. ∎
8. Del Pezzo surfaces of degree 33
Let $S$ be a rational Del Pezzo surface of degree $3$ defined over $\mathbb{Q}$, we want to find the biggest automorphism group $G\subseteq\operatorname{Aut}_{\mathbb{Q}}(S)$. The classification of all automorphism groups of $S$ over $\mathbb{C}$ has been studied in [9, Theorem 9.5.6], and they explained that the Del Pezzo surface of degree $3$ is isomorphic to a one contained in one of the $11$ families of smooth cubic surfaces given in the list.
@@ 8. Del Pezzo surfaces of degree 33
In Theorem 6.1, since M(4,2)=7,M(4,3)=2,M(4,5)=1,M(4,2)=7,~M(4,3)=2,~M(4,5)=1, and $M(4,p)=0$ for all primes $p>5$, then we have
@@
Let $G$ be a finite subgroup of $\operatorname{PGL}_{n}(\mathbb{Q})$ and let $F\in\mathbb{Q}[x_{0},\ldots,x_{n-1}]$ be an irreducible homogeneous polynomial such that $G$ preserves $F$. This means that for each $g$ in $G$ and each matrix $M$ of $\mathrm{GL}_{n}(\mathbb{Q})$ such that g is the class of $M$, the polynomial $F\circ M$ (obtained by acting on the coordinates linearly) is a multiple of $F$. Then, if $d=deg(F)$ and $n$ are coprime, there exists a finite subgroup of $\mathrm{GL}_{n}(\mathbb{Q})$ that surjects to $G$, is twice bigger, and contains Id-Id.
@@
Let $g$ be any element of $G$ and let $M$ be the corresponding matrix. We have $F\circ M=\lambda\cdot F$ for some $\lambda\in\mathbb{Q}$. Let $m$ be the order of $g$. Then, $M^{m}$ is a multiple of the identity, which gives $\mu\in\mathbb{Q}$ such that $M^{m}=\mu\cdot Id$. We write $a=det(M)$. As $M^{m}=\mu\cdot Id$, we apply the determinant on both sides and obtain $a^{m}=\mu^{n}$. Moreover, $F\circ M=\lambda\cdot F$, so $F\circ(M^{i})=\lambda^{i}\cdot F$ for each $i\geq 1$. In particular, $F\circ(M^{m})=\lambda^{m}\cdot F$. As $M^{m}=\mu\cdot Id$, we find $F\circ(\mu\cdot Id)=F\cdot\mu^{d}$ (recall that $d$ is the degree of $F$), so $\mu^{d}=\lambda^{m}$. By Lemma 6.3, there exist $c\in\mathbb{Q}^{*}$ such that $\mu=c^{m}$. Replacing $M$ with $M/c$ we do not change the class $g$ in $G$, but may assume that $M^{m}=Id$. This implies that $\mathrm{det}(M)^{m}=1$, so $\mathrm{det}(M)\in\{-1,1\}$. We can then lift any element of $g\in G$ to an element of $\mathrm{GL}_{n}(\mathbb{Q})$ that has determinant $1$ or 1-1. This gives exactly two possibilities for the lift, namely $M$ or M-M. The set of such matrices makes a finite group twice bigger than $G$. ∎
@@
Let $S$ be a rational Del Pezzo surface defined over $\mathbb{Q}$. Since $\operatorname{Aut}_{\mathbb{Q}}(S)\subseteq\operatorname{Aut}_{\mathbb{C}}(S)$, by [9, Theorem 9.5.6], we only need to study the automorphism groups $G\subset\operatorname{Aut}_{\mathbb{C}}(S)$ such that $\mid G\mid>120$, hence we have only one type to study. If the surface is isomorphic to the Fermat surface $V(t_{0}^{3}+t_{1}^{3}+t_{2}^{3}+t_{3}^{3})$, then $\operatorname{Aut}_{\mathbb{C}}(S)\cong(\mathbb{Z}/3\mathbb{Z})^{3}\times\mathfrak{S}_{4}$ where $\mathfrak{S}_{4}$ is the group of permutations, then $\mid G\mid$ has to divide $648=2^{3}\cdot 3^{4}$. Since $G\subseteq\operatorname{PGL}_{4}(\mathbb{Q})$, then by Remark 8.3, $\mid G\mid$ has to divide $2^{6}\cdot 3^{2}\cdot 5$, hence $\mid G\mid\leq 2^{3}\cdot 3^{2}=48$. ∎
@@
This surface $X$ contains $2$ disjoint (1)(-1)-curves defined over $\mathbb{Q}$ , given by
[…]
By contracting these two (1)(-1)-curves, there is thus a contraction morphism $\eta:X\rightarrow S_{5}$, where $S_{5}$ is a Del Pezzo surface of degree $5$ with at least $2$ rational points, so the surface $S_{5}$ is a rational surface. Indeed, this follows from [17, Theorem 3.15]. Furthermore, the group $G\cong\mathfrak{S}_{5}$ has order $120$.